# EE3P3D: Event-based Estimation of Periodic Phenomena Frequency using 3D Correlation

### v1

@@ 4.3 Measuring frequency of periodic light flashes
The following section presents experiments with frequency estimation methods for estimating high-frequency periodic flicker.
4.3.1 Flashing LED
@@
A simple circuit with a light-emitting diode (see our supplementary) was used for this experiment. A software oscilloscope on the Raspberry Pi Pico controller [5] allowed precise control of the flashing frequency and duty cycle. The LED was set to flash exactly at $2000$ Hz with a $50\%$ duty cycle ($µ250\micro$s of turn-on time). Note that this frequency is substantially higher than those presented in other experiments in this work or related publications (see Sec. 2). The sequence used for this experiment was characterised by near-synchronous event generation across most pixels.
This experiment is the only other experiment (after the ’Felt disc with a high-contrast line’, Sec. 4.2.1) in which all methods achieved near-perfect results (relative error $<0.8\%$) with the Frequency-cam performing the worst (see Tab. 4) The EB-ASM and our proposed method achieved the lowest relative error ($<0.001$%).
4.3.2 Refreshing mobile phone screen
The sequence used for this experiment captured the flicker of a mobile phone screen that displayed a white colour at maximum brightness. A top-to-bottom pattern was visible when inspecting the event data, likely due to Pulse Width Modulation (PWM) for backlight control (see our supplementary). The GT frequency was estimated manually, using the EE3P method and confirmed by findings in a study on the display of this phone [^8] Available at https://bit.ly/notebookcheck-SamsungGalaxyS21UltraReview..
@@ 4.3.2 Refreshing mobile phone screen
Although the Simple baseline method failed, all other methods achieved a relative error lower than $0.8\%$. The limitation of the proposed method is that it requires the aggregation of events to windows of length $t_{\text{quant}}$ (see Sec. 3.3). This restricts both the maximal measurable frequency and the theoretical limits of precision of the method. To illustrate, in this experiment, the ground truth period is $\frac{10^{6}}{240}=4166,\bar{6}$ microseconds. Therefore, the closest estimate of the period ($4200$ microseconds) results in an estimated frequency of $238.1$ Hz and a relative error of $0.792\%$. The Prophesee method and the FFT baseline method achieved the lowest relative errors ($<0.001$%, see Tab. 4).
4.4 Measuring vibration frequency
This section presents two experiments on measuring vibrations of a speaker diaphragm and a motor.
4.4.1 Speaker diaphragm
@@
In this experiment, we captured and analysed a sequence capturing a vibrating diaphragm of a speaker with two large low-frequency drivers. An Android application [^9] Application available at https://bit.ly/Frequency-Sound-Generator-APK. precisely controlled the emitted sound (musical note $G_{2}$, $98$ Hz). The event camera captured the vibrating diaphragm of a speaker at approximately 30° angle. The most prominent features in the event stream originate from the vibrating edges of the diaphragm and the manufacturer’s logo located in its centre (see our supplementary).
The proposed method was the only method that achieved a relative error lower than $0.1\%$. The FFT baseline and Prophesee methods achieved relative error under $0.2\%$, followed by EB-ASM ($1.02\%$) and Frequency-cam ($2.8\%$). The Simple baseline method deviated the most from the reference frequency ($78.571\%$ relative error); see Tab. 4.
4.4.2 Vibrating motor
@@
This experiment analysed a sequence from Prophesee’s dataset capturing a vibrating motor. The most reliable visual features were the horizontal cooling fins of the motor, which periodically moved in a vertical direction (see our supplementary).
The EB-ASM and the Simple baseline method deviated significantly from the GT frequency (relative error $>30\%$, see Tab. 4). Because of an older file encoding, the Frequency-cam method did not provide any measurements. The FFT baseline, Prophesee’s and our proposed method achieved a relative error lower than $0.001$%.
@@ 4.5 Measuring movement frequency
This section presents the application of frequency estimation methods for measuring the speed of moving objects. We present two experiments with a moving bicycle chain in which the frequency at which individual chain links pass is estimated.
4.5.1 Bike chain from side view
The sequence used for this experiment captured the bicycle chain from a close-up side view (see our supplementary).
The FFT baseline method correctly estimated the frequency in pixels that capture the upper and lower chain link sections, where events are periodically generated per chain link passage. However, the estimated frequency in most pixels was incorrect ($1$ Hz, see Tab. 3). The Prophesee’s method deviated by $\approx 18\%$ from the ground truth, being the second most accurate. Note that this is one of only four experiments in which the method by Prophesse achieved a relative error greater than $0.21\%$. Our proposed method achieved the best performance, being the only method to produce meaningful results (see Tab. 3).
4.5.2 Bike chain from top view
In the final experiment, we captured the same bike chain from a top view (see our supplementary).
The Frequency-cam method achieved the second-best result with a $1.3\%$ relative error, with the proposed method achieving the lowest absolute error of $0.1$ Hz. The Simple baseline, Prophesee’s, and EB-ASM methods showed significant deviations from the GT frequency (relative errors $34\%$$6111\%$).
@@ 4.6 Discussion
We analysed twelve sequences capturing various periodic phenomena using six different methods. The experiments where the target lacked high-contrast features were shown to be the most challenging. The worst performing methods were the Simple baseline method (see Sec. 3.1) and the EB-ASM method [2] achieving MRE of $4175.511\%$ and $6848.011\%$ respectively (see Tab. 3,4). The third lowest mean relative error across all experiments ($32.532\%$) was achieved by the Vibration Estimation method by Prophesee. The Frequency-cam method achieved the second lowest mean relative error (MRE) of $14.083\%$, which is significantly higher than the MRE achieved by the proposed method ($0.1\%$).
Note that the Frequency-cam method could not analyse three of the twelve sequences we used in this work, which originated from the public dataset by Prophesee due to an older file encoding type. All other methods were able to analyse all sequences.
Method GT Simple base. FFT base. Prophesee Vibr. Est. EB-ASM Freq-cam EE3P3D (ours)
1 Felt disc with $20.0$ $\mathbf{20.0}$ $\mathbf{20.0}$ $\mathbf{20.0}$ $\mathbf{20.0}$ $\mathbf{20.0}$ $\mathbf{20.0}$
high contrast line
2 Velcro disc (front) $21.1$ $24.2$ $21.0$ $\mathbf{21.1}$ $6250.0$ $21.2$ $\mathbf{21.1}$
3 Velcro disc (side) $26.3$ $26.4$ $53.0$ $\mathbf{26.3}$ $2439.0$ $\mathbf{26.3}$ $\mathbf{26.3}$
4 High contrast dot $19.6$ $7462.7$ $59.0$ $19.5$ $\mathbf{19.6}$ N/A $\mathbf{19.6}$
5 Fidget spinner $4.7$ $451.2$ $5.0$ $19.1$ $1$.11111.11111.11111.1 N/A $\mathbf{4.7}$
6 Whirling spider $3.2$ $35.1$ $13.0$ $4.2$ $298.6$ $5.3$ $\mathbf{3.2}$
7 Flashing LED $2000.0$ $1996.0$ $1999.4$ $2000.1$ $\mathbf{2000.0}$ $1984.1$ $\mathbf{2000.0}$
8 Refreshing mobile phone screen $240.0$ $2770.1$ $\mathbf{240.0}$ $\mathbf{240.0}$ $239.8$ $239.9$ $238.1$
9 Speaker diaphragm $98.0$ $21.0$ $98.1$ $97.8$ $99.0$ $95.3$ $\mathbf{98.0}$
10 Vibrating motor $40.0$ $52.2$ $\mathbf{40.0}$ $\mathbf{40.0}$ $370.4$ N/A $\mathbf{40.0}$
11 Bike chain (side) $28.7$ $123.8$ $1.0$ $33.9$ $1$.11111.11111.11111.1 $44.7$ $\mathbf{28.7}$
12 Bike chain (top) $23.0$ $51.6$ $22.0$ $30.8$ $1428.6$ $22.7$ $\mathbf{23.1}$
Table 3: Frequency estimates (Hz) on the 12 sequence dataset. Methods: Simple baseline (Sec. 3.1), FFT baseline (Sec. 3.2), Prophesse Vibration Estimation [14], EB-ASM [2], Frequency-cam [13] and EE3P3D (Sec. 3.3).
Method Simple base. FFT base. Prophesee Vibr. Est. EB-ASM Freq-cam EE3P3D (ours)
Felt disc with high contrast line $<\epsilon$ $<\epsilon$ $<\epsilon$ $<\epsilon$ $<\epsilon$ $<\epsilon$
Velcro disc (front) $14.6$ $0.5$ $<\epsilon$ $29520.9$ $0.5$ $<\epsilon$
Velcro disc (side) $0.4$ $101.5$ $<\epsilon$ $9173.8$ $<\epsilon$ $<\epsilon$
High contrast dot $37975.0$ $201.0$ $0.5$ $<\epsilon$ N/A $<\epsilon$
Fidget spinner $9500.0$ $6.3$ $306.3$ $23540.4$ N/A $<\epsilon$
Whirling spider $996.9$ $306.3$ $31.3$ $9231.3$ $65.6$ $<\epsilon$
Flashing LED $0.2$ $0.03$ $0.005$ $<\epsilon$ $0.7$ $<\epsilon$
Refreshing mobile phone screen $1054.2$ $<\epsilon$ $<\epsilon$ $0.08$ $0.04$ $0.8$
Speaker diaphragm $78.6$ $0.1$ $0.2$ $1.0$ $2.8$ $<\epsilon$
Vibrating motor $30.5$ $<\epsilon$ $<\epsilon$ $826.0$ N/A $<\epsilon$
Bike chain (side) $331.4$ $96.5$ $18.1$ $3771.4$ $55.8$ $<\epsilon$
Bike chain (top) $124.3$ $4.3$ $33.9$ $6111.3$ $1.3$ $0.4$
Mean relative error $4175.5$ $59.7$ $32.5$ $6848.0$ $14.0$ $0.1$
Max relative error $37975.0$ $306.2$ $306.3$ $29520.9$ $65.6$ $0.8$
Table 4: Relative errors (%) on the 12 sequence dataset. Methods: Simple baseline (Sec. 3.1), FFT baseline (Sec. 3.2), Prophesse Vibration Estimation [14], EB-ASM [2], Frequency-cam [13] and EE3P3D (Sec. 3.3), $\epsilon=10^{-3}$.
@@ 5 Conclusion
In this work, a novel contactless method for measuring the frequency of periodic phenomena is presented. The method makes only a single assumption that the observed object periodically returns to a known state or position, producing a similar set of events in the event stream.
The proposed method, EE3P3D, is evaluated on twelve sequences that represent a wide range of periodic phenomena (light flickering, object vibration, rotational speed, and periodic movement), with targets lacking prominent features, being behind transparent materials, and having centrosymmetric shapes, achieving the mean relative error of $0.1\%$. The method works in non-frontal camera placement, measuring frequencies as high as $2000$ Hz (equivalent to 120 000120000120\,000120 000 RPM) even when a significant amount of noise is present.
The dataset and codes are publicly available on GitHub.
@@ Appendix 0.A Experimental Setup
For the ’Felt disc with a high-contrast line’ experiment (see Sec. 4.2.1), we captured a power drill spinning a white felt disc marked with a black line, see Fig. 0.A.5.
@@
Figure 0.A.5: Felt disc with a high-contrast line (see Sec. 4.2.1): (a) physical setup; (b) events from a 250-millisecond window visualised in spatio-temporal space; (c) aggregated events captured by the event camera ($1280\times 720$px) within a time window of length equal to $\frac{8}{10}$ of the period of the observed phenomenon, highlighted region of interest ($655\times 655$px) shown in (d). Positive events are represented by white colour, and negative events are bright blue.
For the ’Fronto-parallel velcro disc’ experiment (see Sec. 4.2.2), a drill spun the disc at $1266$ RPM with the event camera facing it directly, while a laser tachometer measured speed from the side of the disc where a reflective sticker was attached (Fig. 0.A.6a).
@@
Figure 0.A.6: Fronto-parallel velcro disc (see Sec. 4.2.2): (a) physical setup and target detail; (b) events from a 250-millisecond window visualised in spatio-temporal space; (c) aggregated events captured by the event camera ($1280\times 720$px) within a time window of length equal to the period of the observed phenomenon, highlighted region of interest ($713\times 713$px) shown in (d). Positive events are represented by white colour, and negative events are bright blue.
In the next experiment (’Velcro disc (side)’, see Sec. 4.2.3), the velcro disc was captured by the event camera through a glass sheet at $45\degree$ camera angle (see Fig. 4a).
For the following experiment (’High contrast dot’), a sequence from Prophesee’s dataset was used, capturing an orbiting dark dot (see Sec. 4.2.4). The reference photo was generated using Prophesee’s Event to Video method (see Fig. 0.A.7a).
@@
Figure 0.A.7: High-contrast dot (see Sec. 4.2.4): (a) target visualisation; (b) events from a 250-millisecond window visualised in spatio-temporal space; (c) aggregated events captured by the event camera ($640\times 480$px) within a time window of length equal to $\frac{8}{10}$ of the period of the observed phenomenon, highlighted region of interest ($263\times 263$px) shown in (d). Positive events are represented by white colour, and negative events are bright blue.
This experiment (see Sec. 4.2.5) used event data from Prophesee’s dataset, capturing a rotating three-blade hand fidget spinner on a flat surface, see Fig. 0.A.8.
@@
Figure 0.A.8: Hand fidget spinner (see Sec. 4.2.5): (a) target visualisation; (b) events from a 250-millisecond window visualised in spatio-temporal space; (c) aggregated events captured by the event camera ($640\times 480$px) within a time window of length equal to the period of the observed phenomenon, highlighted region of interest ($239\times 239$px) shown in (d). Positive events are represented by white colour, and negative events are bright blue.
The sequence used for the next experiment captured the whirling behaviour for evading predators used by the Pholcus phalangioides spider (see Sec. 4.2.6). A second rotating pattern is visible, caused by the shadow of the spider (see Fig. 0.A.9d,e).
@@
Figure 0.A.9: Whirling Pholcus phalangioides (see Sec. 4.2.6): (a) physical setup; (b) a close-up photo of the target; (d) aggregated events captured by the event camera ($1280\times 720$px) within a time window of length equal to the period of the observed phenomenon, highlighted region of interest ($221\times 221$px) shown in (c); (e) events from a 250-millisecond window visualised in spatio-temporal space. Positive events are represented by white colour, and negative events are bright blue.
For the LED experiment (see Sec. 4.3.1), a circuit with a light-emitting diode (see Fig. 0.A.10a) was used.
@@
Figure 0.A.10: Flashing LED (see Sec. 4.3.1): (a) physical setup and target detail; (b) events from a 250-millisecond window visualised in spatio-temporal space; (c) aggregated events captured by the event camera ($1280\times 720$px) within a time window of length equal to the period of the observed phenomenon, highlighted region of interest ($126\times 126$px) shown in (d). Positive events are represented by white colour, and negative events are bright blue.
The sequence used for the ’Refreshing mobile phone screen’ experiment (see Sec. 4.3.2) captured the flicker of a mobile phone screen displaying a white colour at maximum brightness. A moving top-to-bottom pattern was visible when inspecting the event data (see Fig. 0.A.11b).
@@
Figure 0.A.11: Refreshing mobile phone screen (see Sec. 4.3.2): (a) physical setup; (b) events from a 250-millisecond window visualised in spatio-temporal space; (c) aggregated events captured within hardware-level region of interest and a time window equal to $\frac{8}{10}$ of the period of the observed phenomenon, highlighted region of interest ($54\times 54$px) shown in (d). Positive events are represented by white colour, and negative events are bright blue.
In the following experiment (see Sec. 4.4.1), we captured a vibrating diaphragm of a speaker with two large low-frequency drivers, see Fig. 0.A.12a.
@@
Figure 0.A.12: Speaker diaphragm (see Sec. 4.4.1): (a) physical setup; (b) events from a 250-millisecond window visualised in spatio-temporal space; (c) aggregated events captured by the event camera ($1280\times 720$px) within a time window of length equal to the period of the observed phenomenon, highlighted region of interest ($631\times 631$px) shown in (d). Positive events are represented by white colour, and negative events are bright blue.
For the ’Vibrating motor’ experiment (see Sec. 4.4.2) we used a sequence from Prophesee’s dataset capturing a motor vibrating vertically at $40$ Hz; see Fig. 0.A.13.
@@
Figure 0.A.13: Vibrating motor (see Sec. 4.4.2): (a) target visualisation; (b) events from a 250-millisecond window visualised in spatio-temporal space; (c) aggregated events captured by the event camera ($640\times 480$px) within a time window of length equal to the period of the observed phenomenon, highlighted region of interest ($148\times 148$px) shown in (d). Positive events are represented by white colour, and negative events are bright blue.
The sequence used for this experiment (see Sec. 4.5.1) captured a bicycle chain from a close-up side view (see Fig. 0.A.14a).
@@
Figure 0.A.14: Bike chain from side view (see Sec. 4.5.1): (a) physical setup; (b) events from a 250-millisecond window visualised in spatio-temporal space; (c) aggregated events captured within a hardware-level region of interest and a time window equal to the period of the observed phenomenon, highlighted region of interest ($204\times 204$px) shown in (d). Positive events are represented by white colour, and negative events are bright blue.
The sequence for the final experiment (see Sec. 4.5.2) captured the same bike chain from a top view perspective (see Fig. 0.A.15a).
@@
Figure 0.A.15: Bike chain from a top view (see Sec. 4.5.2): (a) physical setup; (b) events from a 250-millisecond window visualised in spatio-temporal space; (c) aggregated events captured by the event camera ($1280\times 720$px) within a time window of lenght equal to the period of the observed phenomenon, highlighted region of interest ($180\times 180$px) shown in (d). Positive events are represented by white colour, and negative events are bright blue.

### v2

@@ 5.3 Measuring frequency of periodic light flashes
Flashing LED
@@
A simple circuit with a light-emitting diode was used for this experiment. A software oscilloscope on the Raspberry Pi Pico controller [5] allowed precise control of the flashing frequency and duty cycle. The LED was set to flash exactly at $2000$ Hz with a $50\%$ duty cycle ($µ250\micro$s of turn-on time). Note that this frequency is substantially higher than those presented in other experiments in this work or related publications (see Sec. 2). The sequence used for this experiment was characterised by near-synchronous event generation across most pixels. This experiment is the only other experiment (after the ’Felt disc with a high-contrast line’, Sec. 5.2) in which all methods achieved near-perfect results (relative error $<0.8\%$) with the Frequency-cam performing the worst (see Tab. 3) The EB-ASM and our proposed method achieved the lowest relative error ($<0.001$%).
Refreshing screen
@@
The sequence used for this experiment captured the flicker of a mobile phone screen that displayed a white colour at maximum brightness. A top-to-bottom pattern was visible when inspecting the event data, likely due to Pulse Width Modulation (PWM) for backlight control. The GT rate was estimated manually, using the EE3P method and confirmed by findings in a study on the display of this phone [^7] Available at https://bit.ly/notebookcheck-SamsungGalaxyS21UltraReview.. Although the Simple baseline method failed, all other methods achieved a relative error lower than $0.8\%$. The limitation of the proposed method is that it requires the aggregation of events to windows of length $t_{\text{quant}}$ (see Sec. 4). This restricts both the maximal measurable frequency and the theoretical limits of precision of the method. To illustrate, in this experiment, the ground truth period is $\frac{10^{6}}{240}=4166,\bar{6}$ microseconds. Therefore, the closest estimate of the period ($4200$ microseconds) results in an estimated rate of $238.1$ Hz and a relative error of $0.792\%$. The Prophesee method and the FFT baseline method achieved the lowest relative errors ($<0.001$%, see Tab. 3).
Frequency Estimates (Hz) Relative Errors (%)
Method GT Simple base. FFT base. Prophesee Vibr. Est. EB-ASM Freq-cam Ours Simple base. FFT base. Prophesee Vibr. Est. EB-ASM Freq-cam Ours
1 Felt disc with high contrast line 20.0 20.0 20.0 20.0 20.0 20.0 20.0 $<\epsilon$ $<\epsilon$ $<\epsilon$ $<\epsilon$ $<\epsilon$ $<\epsilon$
2 Velcro disc (front) 21.1 24.2 21.0 21.1 6250.0 21.2 21.1 14.6 0.5 $<\epsilon$ 29520.9 0.5 $<\epsilon$
3 Velcro disc (side) 26.3 26.4 53.0 26.3 2439.0 26.3 26.3 0.4 101.5 $<\epsilon$ 9173.8 $<\epsilon$ $<\epsilon$
4 High contrast dot 19.6 7462.7 59.0 19.5 19.6 N/A 19.6 37975.0 201.0 0.5 $<\epsilon$ N/A $<\epsilon$
5 Fidget spinner 4.7 451.2 5.0 19.1 $1$.1 N/A 4.7 9500.0 6.3 306.3 23540.4 N/A $<\epsilon$
6 Whirling spider 3.2 35.1 13.0 4.2 298.6 5.3 3.2 996.9 306.3 31.3 9231.3 65.6 $<\epsilon$
7 Flashing LED 2000.0 1996.0 1999.4 2000.1 2000.0 1984.1 2000.0 0.2 0.03 0.005 $<\epsilon$ 0.7 $<\epsilon$
8 Refreshing screen 240.0 2770.1 240.0 240.0 239.8 239.9 238.1 1054.2 $<\epsilon$ $<\epsilon$ 0.08 0.04 0.8
9 Speaker diaphragm 98.0 21.0 98.1 97.8 99.0 95.3 98.0 78.6 0.1 0.2 1.0 2.8 $<\epsilon$
10 Vibrating motor 40.0 52.2 40.0 40.0 370.4 N/A 40.0 30.5 $<\epsilon$ $<\epsilon$ 826.0 N/A $<\epsilon$
11 Bike chain (side) 28.7 123.8 1.0 33.9 $1$.1 44.7 28.7 331.4 96.5 18.1 3771.4 55.8 $<\epsilon$
12 Bike chain (top) 23.0 51.6 22.0 30.8 1428.6 22.7 23.1 124.3 4.3 33.9 6111.3 1.3 0.4
Table 3: Frequency estimates (Hz) and relative errors (%) on the 12 sequence dataset. Methods: Simple baseline (Sec. 3), FFT baseline (Sec. 3), Prophesee Vibration Estimation [14], EB-ASM [1], Frequency-cam [13] and EEPPR (Sec. 4).
5.4 Measuring vibration frequency of a speaker and a motor
Speaker diaphragm
@@ 5.4 Measuring vibration frequency of a speaker and a motor
In this experiment, we captured and analysed a sequence capturing a vibrating diaphragm of a speaker with two large low-frequency drivers. An Android application [^8] Application available at https://bit.ly/Frequency-Sound-Generator-APK. precisely controlled the emitted sound (musical note $G_{2}$, $98$ Hz). The event camera captured the vibrating diaphragm of a speaker at approximately 30° angle. The most prominent features in the event stream originate from the vibrating edges of the diaphragm and the manufacturer’s logo located in its centre. The proposed method was the only method that achieved a relative error lower than $0.1\%$. The FFT baseline and Prophesee methods achieved relative error under $0.2\%$, followed by EB-ASM ($1.02\%$) and Frequency-cam ($2.8\%$). The Simple baseline method deviated the most from the reference frequency ($78.571\%$ relative error); see Tab. 3.
Vibrating motor
@@
This experiment analysed a sequence from Prophesee’s dataset capturing a vibrating motor. The most reliable visual features were the horizontal cooling fins of the motor, which periodically moved in a vertical direction. The EB-ASM and the Simple baseline method deviated significantly from the GT frequency (relative error $>30\%$, see Tab. 3). Because of an older file encoding, the Frequency-cam method did not provide any measurements. The FFT baseline, Prophesee’s and our proposed method achieved a relative error lower than $0.001$%.
@@ 5.5 Measuring movement frequency
Bike chain from side view
The sequence used for this experiment captured the bicycle chain from a close-up side view. The FFT baseline method correctly estimated the frequency in pixels that capture the upper and lower chain link sections, where events are periodically generated per chain link passage. However, the estimated frequency in most pixels was incorrect ($1$ Hz, see Tab. 3). The Prophesee method deviated by $\approx 18\%$ from the ground truth, being the second most accurate; this is one of only four experiments where the method by Prophesse had a relative error greater than $0.21\%$. EEPPR achieved the best performance, being the only method to produce meaningful results.
Bike chain from top view
In the final experiment, we captured the same bike chain from a top view. The Frequency-cam method achieved the second-best result with a $1.3\%$ relative error, with the proposed method achieving the lowest absolute error of $0.1$ Hz. The Simple baseline, Prophesee’s, and EB-ASM methods showed significant deviations from the GT frequency (relative errors $34\%$$6111\%$).
@@ 5.6 Discussion
We analysed twelve sequences capturing various periodic phenomena using six different methods. The experiments where the target lacked high-contrast features were shown to be the most challenging. The worst performing methods were the Simple baseline method and the EB-ASM method [1] achieving MRE of $4175.511\%$ and $6848.011\%$ respectively (see Tab. 3). The third lowest mean relative error across all experiments ($32.532\%$) was achieved by the Vibration Estimation method by Prophesee. The Frequency-cam method achieved the second lowest mean relative error (MRE) of $14.083\%$, which is significantly higher than the MRE achieved by the proposed method ($0.1\%$). Note that the Frequency-cam method could not analyse three of the twelve sequences we used in this work, which originated from the public dataset by Prophesee due to an older file encoding type. All other methods were able to analyse all sequences.
@@ 6 Conclusion
A novel contactless method for measuring the rate of periodic phenomena has been presented. The method makes a single assumption that the observed object periodically returns to a known state or position, producing a similar set of events in the event stream. The method, EEPPR, was evaluated on a new dataset of twelve sequences that represent a wide range of periodic phenomena (light flickering, object vibration, rotational speed, and periodic movement), with targets lacking prominent features, being behind transparent materials, and having centrosymmetric shapes, achieving the mean relative error of $0.1\%$. The method works in non-frontal camera placement, measuring frequencies as high as $2000$ Hz (equivalent to 120 000120000120\,000120 000 RPM) even when a significant amount of noise is present.
Limitations
The method assumes the location where the period phenomenon is occurring is static. Handling of tracking and the template update is future work.
@@ Appendix 0.A Experimental Setup
For the ’Felt disc with a high-contrast line’ experiment (see Sec. 5.2), we captured a power drill spinning a white felt disc marked with a black line, see Fig. 0.A.5.
@@
Figure 0.A.5: Felt disc with a high-contrast line (see Sec. 5.2): (a) physical setup; (b) events from a 250-millisecond window visualised in spatio-temporal space; (c) aggregated events captured by the event camera ($1280\times 720$px) within a time window of length equal to $\frac{8}{10}$ of the period of the observed phenomenon, highlighted region of interest ($655\times 655$px) shown in (d). Positive events are represented by white colour, and negative events are bright blue.
For the ’Fronto-parallel velcro disc’ experiment (see Sec. 5.2), a drill spun the disc at $1266$ RPM with the event camera facing it directly, while a laser tachometer measured speed from the side of the disc where a reflective sticker was attached (Fig. 0.A.6a).
@@
Figure 0.A.6: Fronto-parallel velcro disc (see Sec. 5.2): (a) physical setup and target detail; (b) events from a 250-millisecond window visualised in spatio-temporal space; (c) aggregated events captured by the event camera ($1280\times 720$px) within a time window of length equal to the period of the observed phenomenon, highlighted region of interest ($713\times 713$px) shown in (d). Positive events are represented by white colour, and negative events are bright blue.
In the next experiment (’Velcro disc (side)’, see Sec. 5.2), the velcro disc was captured by the event camera through a glass sheet at $45\degree$ camera angle (see Fig. 4a).
For the following experiment (’High contrast dot’), a sequence from Prophesee’s dataset was used, capturing an orbiting dark dot (see Sec. 5.2). The reference photo was generated using Prophesee’s Event to Video method (see Fig. 0.A.7a).
@@
Figure 0.A.7: High-contrast dot (see Sec. 5.2): (a) target visualisation; (b) events from a 250-millisecond window visualised in spatio-temporal space; (c) aggregated events captured by the event camera ($640\times 480$px) within a time window of length equal to $\frac{8}{10}$ of the period of the observed phenomenon, highlighted region of interest ($263\times 263$px) shown in (d). Positive events are represented by white colour, and negative events are bright blue.
This experiment (see Sec. 5.2) used event data from Prophesee’s dataset, capturing a rotating three-blade hand fidget spinner on a flat surface, see Fig. 0.A.8.
@@
Figure 0.A.8: Hand fidget spinner (see Sec. 5.2): (a) target visualisation; (b) events from a 250-millisecond window visualised in spatio-temporal space; (c) aggregated events captured by the event camera ($640\times 480$px) within a time window of length equal to the period of the observed phenomenon, highlighted region of interest ($239\times 239$px) shown in (d). Positive events are represented by white colour, and negative events are bright blue.
The sequence used for the next experiment captured the whirling behaviour for evading predators used by the Pholcus phalangioides spider (see Sec. 5.2). A second rotating pattern is visible, caused by the shadow of the spider (see Fig. 0.A.9d,e).
@@
Figure 0.A.9: Whirling Pholcus phalangioides (see Sec. 5.2): (a) physical setup; (b) a close-up photo of the target; (d) aggregated events captured by the event camera ($1280\times 720$px) within a time window of length equal to the period of the observed phenomenon, highlighted region of interest ($221\times 221$px) shown in (c); (e) events from a 250-millisecond window visualised in spatio-temporal space. Positive events are represented by white colour, and negative events are bright blue.
For the LED experiment (see Sec. 5.3), a circuit with a light-emitting diode (see Fig. 0.A.10a) was used.
@@
Figure 0.A.10: Flashing LED (see Sec. 5.3): (a) physical setup and target detail; (b) events from a 250-millisecond window visualised in spatio-temporal space; (c) aggregated events captured by the event camera ($1280\times 720$px) within a time window of length equal to the period of the observed phenomenon, highlighted region of interest ($126\times 126$px) shown in (d). Positive events are represented by white colour, and negative events are bright blue.
The sequence used for the ’Refreshing mobile phone screen’ experiment (see Sec. 5.3) captured the flicker of a mobile phone screen displaying a white colour at maximum brightness. A moving top-to-bottom pattern was visible when inspecting the event data (see Fig. 0.A.11b).
@@
Figure 0.A.11: Refreshing mobile phone screen (see Sec. 5.3): (a) physical setup; (b) events from a 250-millisecond window visualised in spatio-temporal space; (c) aggregated events captured within hardware-level region of interest and a time window equal to $\frac{8}{10}$ of the period of the observed phenomenon, highlighted region of interest ($54\times 54$px) shown in (d). Positive events are represented by white colour, and negative events are bright blue.
In the following experiment (see Sec. 5.4), we captured a vibrating diaphragm of a speaker with two large low-frequency drivers, see Fig. 0.A.12a.
@@
Figure 0.A.12: Speaker diaphragm (see Sec. 5.4): (a) physical setup; (b) events from a 250-millisecond window visualised in spatio-temporal space; (c) aggregated events captured by the event camera ($1280\times 720$px) within a time window of length equal to the period of the observed phenomenon, highlighted region of interest ($631\times 631$px) shown in (d). Positive events are represented by white colour, and negative events are bright blue.
For the ’Vibrating motor’ experiment (see Sec. 5.4) we used a sequence from Prophesee’s dataset capturing a motor vibrating vertically at $40$ Hz; see Fig. 0.A.13.
@@
Figure 0.A.13: Vibrating motor (see Sec. 5.4): (a) target visualisation; (b) events from a 250-millisecond window visualised in spatio-temporal space; (c) aggregated events captured by the event camera ($640\times 480$px) within a time window of length equal to the period of the observed phenomenon, highlighted region of interest ($148\times 148$px) shown in (d). Positive events are represented by white colour, and negative events are bright blue.
The sequence used for this experiment (see Sec. 5.5) captured a bicycle chain from a close-up side view (see Fig. 0.A.14a).
@@
Figure 0.A.14: Bike chain from side view (see Sec. 5.5): (a) physical setup; (b) events from a 250-millisecond window visualised in spatio-temporal space; (c) aggregated events captured within a hardware-level region of interest and a time window equal to the period of the observed phenomenon, highlighted region of interest ($204\times 204$px) shown in (d). Positive events are represented by white colour, and negative events are bright blue.
The sequence for the final experiment (see Sec. 5.5) captured the same bike chain from a top view perspective (see Fig. 0.A.15a).
@@
Figure 0.A.15: Bike chain from a top view (see Sec. 5.5): (a) physical setup; (b) events from a 250-millisecond window visualised in spatio-temporal space; (c) aggregated events captured by the event camera ($1280\times 720$px) within a time window of lenght equal to the period of the observed phenomenon, highlighted region of interest ($180\times 180$px) shown in (d). Positive events are represented by white colour, and negative events are bright blue.